OpenStudy (anonymous):

OK, Express the equation in logarithm form (use the same variable): e^4x+40=200 WHICH I FOUND: LN(200)=4X+40 BUT how do i Solve the equation for x. x =

OpenStudy (anonymous):

should this look like e^(4x+40)=200 or e^(4x)+40=200 ?

OpenStudy (anonymous):

the answer? or the equation?

OpenStudy (anonymous):

the equation

OpenStudy (anonymous):

yes the first one

OpenStudy (anonymous):

then you just solve for x, ln(200) is not a "pretty" number so leave it as the ln its the exact answer instead of trying to round it

OpenStudy (anonymous):

ok but how can i write it out w/ ln(200) as is? i mean 4x+40 part

OpenStudy (anonymous):

because they want me to solve for x

OpenStudy (anonymous):

ln(200) is jsut a number like any other just solve forxlike you would ifit was n=4x+40

OpenStudy (anonymous):

so it would be x= ln(200)-40/4??

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thanx